OpenStudy (anonymous):

I just need someone to check my answers in using matrices to determine the coordinates of matrices.

OpenStudy (anonymous):

http://prntscr.com/8g2uxw

OpenStudy (anonymous):

For both answers I got C. http://prntscr.com/8g2v9u

OpenStudy (lynfran):

for this 1...its A http://prntscr.com/8g2v9u

OpenStudy (anonymous):

Ah, that was my first assumption, but I second guessed myself.

OpenStudy (lynfran):

and for the other its A too

OpenStudy (lynfran):

we use the x axis matrix \[\left[\begin{matrix}1 & 0 \\ 0 & -1\end{matrix}\right]\]times it by each vertex to find solution of its it image

OpenStudy (lynfran):

the 270 degree matrix \[\left[\begin{matrix}0 & 1 \\ -1 & 0\end{matrix}\right]\]we do the same here too

OpenStudy (anonymous):

Thanks! I really appreciate your explanations because matrices have been giving me such a hard time.

OpenStudy (lynfran):

example with vertices V(8,4) and we want its image that was revolve 270 degree we multiply that by its matrix \[\left[\begin{matrix}0 & 1 \\ -1 & 0\end{matrix}\right]*\left(\begin{matrix}8 \\ 4\end{matrix}\right)\]\[\left(\begin{matrix}0*8+(-1)*4 \\ (-1)*8+0*4\end{matrix}\right)\]we get \[\left(\begin{matrix}-4 \\ -8\end{matrix}\right)\]as the image of V

OpenStudy (anonymous):

Ah, I see. Do you think you can help with scale factor with matrices as well?

OpenStudy (lynfran):

ok

OpenStudy (anonymous):

So this is my question. http://prntscr.com/8g302o I got A, but I was just guessing since the others didn't look right, but the points aren't exact, so it's kind of confusing.

OpenStudy (anonymous):

Actually I think it's this one! http://prntscr.com/8g30iv

OpenStudy (lynfran):